OpenStudy (anonymous):

given integral e^(-x^2) from -infinity to + infinity =pi^2 calculate the exact value of integral e^-(x-a)^2/b from -infinity to + infinity

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} e^(-x^2)dx = \sqrt{\pi}\] calculate \[\int\limits_{-\infty}^{\infty}e^-(x-a)^2/b\]

OpenStudy (jamesj):

we clearly the -a has no effect, right?, because you can make a change of variables u = x - a and not introduce any constants. Ok, now, make one more change of variables to get rid of the b in the exponent and what do you get?

OpenStudy (anonymous):

The main question is what does pi have to do with the Gaussian function?

OpenStudy (anonymous):

let me prove this statment

OpenStudy (anonymous):

i know the answer don't rush?

OpenStudy (jamesj):

That is not the main question. The pi is given and it is not asked to proved. I think you are going to confuse here by proving that relationship.

OpenStudy (anonymous):

i am going to prove the whole mathematical concept

OpenStudy (anonymous):

yes i know but the thing is if no one learns how to derive things then they will always forget the answer

OpenStudy (anonymous):

if they don't understand the relationship then is this all about just getting the right answer to score a grade they don;t earn

OpenStudy (anonymous):

i was tending to use the jacobians febnus method to prove the gaussian function

OpenStudy (anonymous):

which uses matrix determinant and double integral, this proof was innovated by laplace

OpenStudy (anonymous):

also polar coordinates

OpenStudy (anonymous):

you can also use erf with limits to prove the concept

OpenStudy (anonymous):

i'll leave this page but if required the proof let me know

OpenStudy (whpalmer4):

@Ldaniel to typeset e^(expr), use e^{expr} For example, \[e^{-(x-a)^2}\] is \ [ e ^ { - (x - a ) ^ 2 } \ ] (without the spaces) By default only the next character after the ^ symbol gets treated as an exponent unless you do that.

OpenStudy (anonymous):

so how do I calculate the exact value? I don't know what to do with "b"

OpenStudy (anonymous):

\[ \int\limits_{-\infty}^{\infty}e{ (x-a )^2}/b\]

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} e ^{-(u)^2}/b\] u=(x-a)

OpenStudy (anonymous):

thanks by the way @whpalmer4

OpenStudy (whpalmer4):

You're welcome. It was all I could do at that hour with the amount of blood in my caffeine stream :-)

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} e^ {(-(x-a)^2)/b} dx\] this this the right integral

OpenStudy (sirm3d):

hint: \[\Large \frac{(x-a)^2}{b}=\left(\frac{x-a}{\sqrt {b}}\right)^2\]

OpenStudy (phi):

to do this problem, use sirm's idea and do a change of variables. let \[ u^2 = \frac{(x-a)^2}{b} \] \[ u = \frac{x}{\sqrt{b}}- \frac{a}{\sqrt{b}} \] we need du in terms of dx (or vice versa). take the derivative of both sides, treating a and b as constants, and u and x as variables: \[ du = \frac{1}{\sqrt{b}} dx \] and, in terms of dx, \[ \sqrt{b} \text{ }du = dx\] we also need the limits for u. when x = -inf, we find \[ u = \frac{-\infty}{\sqrt{b}}- \frac{a}{\sqrt{b}} = -\infty \] similarly the upper limit for u is \( +\infty \) now do the change of variables \[ \int e^{-u^2} \sqrt{b} du = \sqrt{b} \int_{-\infty}^{+\infty} e^{-u^2} du \] of course we could write this as \[ \sqrt{b} \int_{-\infty}^{+\infty} e^{-x^2} dx \]

OpenStudy (anonymous):

\[\sqrt{\pi b}\]

OpenStudy (anonymous):

right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you so much @phi

OpenStudy (phi):

The main thing is learn how to do the change of variables. It is a useful skill.

OpenStudy (anonymous):

yeah